OpenStudy (anonymous):

2. Find the next term in this arithmetic sequence: 2/3, 2, 3 1/3

OpenStudy (jlastino):

You should find the difference between the second term and first term 2- 2/3 = ? Then add that to 3 1/3 to get the missing term

OpenStudy (anonymous):

\[x=n+1 _{1/3}\]

OpenStudy (anonymous):

\[3^{1/3} + 1^{1/3} = 4^{2/3}\]

OpenStudy (anonymous):

\[3^{1/3}-2=1^{2/3}\]